Desiree is twice as old as her cousin now. In 30 years time, Desiree age will be 14 years more than 2/3 of her cousin's age.
a) How old is Desiree now?

Answer by ewatrrr(24785)


Hi
Re: TY (IF 14 was on the left side..it would need to be -14)
Desiree's age: 2x (cousin x)
in 30 years...Desiree age will be 14 years more than 2/3 of her cousin's age
2x + 30 = 2/3(x+30) + 14
2x + 30 = (2/3)x + (2/3)30 + 14
(4/3)x = 20+14-30
(4/3)x = 4
x = 3, cousin's age, Desiree is 6
CHECKING our answer***in 30 years
36 = 2/3(33) + 14 = 22 + 14 = 36
Wish You the Best in your Studies.